the length of a rectangular painting is 3 inches. longer than it's width. If the diagonal is 15 inches long, what are the dimensions of the painting?
Answer by fractalier(6550)

If we call the width, x, then the length is x + 3.
We can apply the Pythagorean Theorem and write
x^2 + (x+3)^2 = 15^2
If you work that out, you'll find that the width
x = 9 inches
and
length x+3 = 12 inches
It's a 9 x 12.